# MMSR: Symbolic Regression is a Multimodal Task

### v4

@@ Appendix C Appendix: Test data in detail
II.2.42 P $=\frac{\kappa(T_{2}-T_{1})A}{d}$ 0.8015
II.3.24 $F_{E}=\frac{P}{4\pi r^{2}}$ 0.9813
II.4.23 $V_{e}=\frac{q}{4\pi\epsilon r}$ 0.9972
II.6.11 $V_{e}=\frac{1}{4\pi\epsilon}\frac{p_{d}\cos\theta}{r^{2}}$ 0.9883
II.6.15a $E_{f}=\frac{3}{4\pi\epsilon}\frac{p_{d}z}{r^{5}}\sqrt{x^{2}+y^{2}}$ 0.9221
II.6.15b $E_{f}=\frac{3}{4\pi\epsilon}\frac{p_{d}}{r^{3}}\cos\theta\sin\theta$ 0.9917
II.8.7 $E=\frac{3}{5}\frac{q^{2}}{4\pi\epsilon d}$ 0.9816
II.8.31 $E_{den}=\frac{\epsilon E_{f}^{2}}{2}$ 1.0
II.10.9 $E_{f}=\frac{\sigma_{den}}{\epsilon}\frac{1}{1+\chi}$ 0.9999
II.11.3 $x=\frac{qE_{f}}{m(\omega_{0}^{2}-\omega^{2})}$ 0.9824
II.11.7 $n=n_{0}(1+\frac{p_{d}E_{f}\cos\theta}{k_{b}T})$ 0.8729
II.11.20 $P_{*}=\frac{n_{\rho}p_{d}^{2}E_{f}}{3k_{b}T}$ 0.7225
II.11.27 $P_{*}=\frac{n\alpha}{1-n\alpha/3}\epsilon E_{f}$ 0.9817
II.11.28 $\theta=1+\frac{n\alpha}{1-(n\alpha/3)}$ 0.9991
II.13.17 $B=\frac{1}{4\pi\epsilon c^{2}}\frac{2I}{r}$ 0.9961
II.13.23 $\rho_{c}=\frac{\rho_{c_{0}}}{\sqrt{1-v^{2}/c^{2}}}$ 0.9622
II.13.34 $j=\frac{\rho_{c_{0}}v}{\sqrt{1-v^{2}/c^{2}}}$ 0.9847
@@
II.21.32 $V_{e}=\frac{q}{4\pi\epsilon r(1-v/c)}$ 0.9915
II.24.17 $k=\sqrt{\frac{\omega^{2}}{c^{2}}-\frac{\pi^{2}}{d^{2}}}$ 0.9872
II.27.16 $F_{E}=\epsilon cE_{f}^{2}$ 0.9917
II.27.18 $E_{den}=\epsilon E_{f}^{2}$ 0.9993
II.34.2a $I=\frac{qv}{2\pi r}$ 0.9916
II.34.2 $\mu_{M}=\frac{qvr}{2}$ 0.9862
II.34.11 $\omega=\frac{g_{\_}qB}{2m}$ 0.9926
II.34.29a $\mu_{M}=\frac{qh}{4\pi m}$ 0.9987
II.34.29b $E=\frac{g_{\_}\mu_{M}BJ_{z}}{\hbar}$ 0.8219
II.35.18 $n=\frac{n_{0}}{\exp(\mu_{m}B/(k_{b}T))+\exp(-\mu_{m}B/(k_{b}T))}$ 0.9512
II.35.21 $M=n_{\rho}\mu_{M}\tanh(\frac{\mu_{M}B}{k_{b}T})$ 0.8199
II.36.38 $f=\frac{\mu_{m}B}{k_{b}T}+\frac{\mu_{m}\alpha M}{\epsilon c^{2}k_{b}T}$ 0.9250
II.37.1 $E=\mu_{M}(1+\chi)B$ 0.9999
@@
II.38.14 $\mu_{S}=\frac{Y}{2(1+\sigma)}$ 0.9999
III.4.32 $n=\frac{1}{e^{\frac{\hbar\omega}{k_{b}T}}-1}$ 0.9628
III.4.33 $E=\frac{\hbar\omega}{e^{\frac{\hbar\omega}{k_{b}T}}-1}$ 0.9973
III.7.38 $\omega=\frac{2\mu_{M}B}{\hbar}$ 0.9826
III.8.54 $p_{\gamma}$ $=\sin(\frac{Et}{\hbar})^{2}$ 0.9822
III.9.52 $p_{\gamma}$ $=\frac{p_{d}E_{f}t}{\hbar}\frac{\sin((\omega-\omega_{0})t/2)^{2}}{((\omega-\omega_{0})t/2)^{2}}$ 0.7024
III.10.19 $E=\mu_{M}\sqrt{B_{x}^{2}+B_{y}^{2}+B_{z}^{2}}$ 0.9948
III.12.43 $L=n\hbar$ 0.9924
III.13.18 $v=\frac{2Ed^{2}k}{\hbar}$ 0.9999
III.14.14 $I=I_{0}(e^{\frac{qV_{e}}{k_{b}T}}-1)$ 0.9914
@@
III.15.14 $m=\frac{\hbar^{2}}{2Ed^{2}}$ 0.9995
III.15.27 $k=\frac{2\pi\alpha}{nd}$ 0.9914
III.17.37 $f=\beta(1+\alpha\cos\theta)$ 0.9988
III.19.51 $E=\frac{-mq^{4}}{2(4\pi\epsilon)^{2}\hbar^{2}}\frac{1}{n^{2}}$ 0.9894
III.21.20 $j=\frac{-\rho_{c_{0}}qA_{vec}}{m}$ 0.7262
Table 6: Tested Feynman Equations, part 2.

### v5

@@ Appendix C Appendix: Test data in detail
II.2.42 P $=\frac{\kappa(T_{2}-T_{1})A}{d}$ 0.8511
II.3.24 $F_{E}=\frac{P}{4\pi r^{2}}$ 0.9910
II.4.23 $V_{e}=\frac{q}{4\pi\epsilon r}$ 0.9892
II.6.11 $V_{e}=\frac{1}{4\pi\epsilon}\frac{p_{d}\cos\theta}{r^{2}}$ 0.9939
II.6.15a $E_{f}=\frac{3}{4\pi\epsilon}\frac{p_{d}z}{r^{5}}\sqrt{x^{2}+y^{2}}$ 0.9650
II.6.15b $E_{f}=\frac{3}{4\pi\epsilon}\frac{p_{d}}{r^{3}}\cos\theta\sin\theta$ 0.9928
II.8.7 $E=\frac{3}{5}\frac{q^{2}}{4\pi\epsilon d}$ 0.9906
II.8.31 $E_{den}=\frac{\epsilon E_{f}^{2}}{2}$ 0.9999
II.10.9 $E_{f}=\frac{\sigma_{den}}{\epsilon}\frac{1}{1+\chi}$ 0.9995
II.11.3 $x=\frac{qE_{f}}{m(\omega_{0}^{2}-\omega^{2})}$ 0.9904
II.11.7 $n=n_{0}(1+\frac{p_{d}E_{f}\cos\theta}{k_{b}T})$ 0.8918
II.11.20 $P_{*}=\frac{n_{\rho}p_{d}^{2}E_{f}}{3k_{b}T}$ 0.8311
II.11.27 $P_{*}=\frac{n\alpha}{1-n\alpha/3}\epsilon E_{f}$ 0.9955
II.11.28 $\theta=1+\frac{n\alpha}{1-(n\alpha/3)}$ 0.9999
II.13.17 $B=\frac{1}{4\pi\epsilon c^{2}}\frac{2I}{r}$ 0.9853
II.13.23 $\rho_{c}=\frac{\rho_{c_{0}}}{\sqrt{1-v^{2}/c^{2}}}$ 0.9723
II.13.34 $j=\frac{\rho_{c_{0}}v}{\sqrt{1-v^{2}/c^{2}}}$ 0.9907
@@
II.21.32 $V_{e}=\frac{q}{4\pi\epsilon r(1-v/c)}$ 0.9983
II.24.17 $k=\sqrt{\frac{\omega^{2}}{c^{2}}-\frac{\pi^{2}}{d^{2}}}$ 0.9952
II.27.16 $F_{E}=\epsilon cE_{f}^{2}$ 0.9991
II.27.18 $E_{den}=\epsilon E_{f}^{2}$ 0.9989
II.34.2a $I=\frac{qv}{2\pi r}$ 0.9938
II.34.2 $\mu_{M}=\frac{qvr}{2}$ 0.9935
II.34.11 $\omega=\frac{g_{\_}qB}{2m}$ 0.9903
II.34.29a $\mu_{M}=\frac{qh}{4\pi m}$ 0.9830
II.34.29b $E=\frac{g_{\_}\mu_{M}BJ_{z}}{\hbar}$ 0.8013
II.35.18 $n=\frac{n_{0}}{\exp(\mu_{m}B/(k_{b}T))+\exp(-\mu_{m}B/(k_{b}T))}$ 0.9244
II.35.21 $M=n_{\rho}\mu_{M}\tanh(\frac{\mu_{M}B}{k_{b}T})$ 0.7822
II.36.38 $f=\frac{\mu_{m}B}{k_{b}T}+\frac{\mu_{m}\alpha M}{\epsilon c^{2}k_{b}T}$ 0.9550
II.37.1 $E=\mu_{M}(1+\chi)B$ 0.9924
@@
II.38.14 $\mu_{S}=\frac{Y}{2(1+\sigma)}$ 0.9994
III.4.32 $n=\frac{1}{e^{\frac{\hbar\omega}{k_{b}T}}-1}$ 0.9322
III.4.33 $E=\frac{\hbar\omega}{e^{\frac{\hbar\omega}{k_{b}T}}-1}$ 0.9976
III.7.38 $\omega=\frac{2\mu_{M}B}{\hbar}$ 0.9922
III.8.54 $p_{\gamma}$ $=\sin(\frac{Et}{\hbar})^{2}$ 0.9723
III.9.52 $p_{\gamma}$ $=\frac{p_{d}E_{f}t}{\hbar}\frac{\sin((\omega-\omega_{0})t/2)^{2}}{((\omega-\omega_{0})t/2)^{2}}$ 0.7323
III.10.19 $E=\mu_{M}\sqrt{B_{x}^{2}+B_{y}^{2}+B_{z}^{2}}$ 0.9838
III.12.43 $L=n\hbar$ 0.9794
III.13.18 $v=\frac{2Ed^{2}k}{\hbar}$ 0.9989
III.14.14 $I=I_{0}(e^{\frac{qV_{e}}{k_{b}T}}-1)$ 0.9825
@@
III.15.14 $m=\frac{\hbar^{2}}{2Ed^{2}}$ 0.9990
III.15.27 $k=\frac{2\pi\alpha}{nd}$ 0.9982
III.17.37 $f=\beta(1+\alpha\cos\theta)$ 0.9413
III.19.51 $E=\frac{-mq^{4}}{2(4\pi\epsilon)^{2}\hbar^{2}}\frac{1}{n^{2}}$ 0.9739
III.21.20 $j=\frac{-\rho_{c_{0}}qA_{vec}}{m}$ 0.8113
Table D.2: Tested Feynman Equations, part 2.